Megan has 3 game scores that happen to be square numbers. The first 2 scores have the same 3 digits the total of the 3 scores is 590. what are the 3 scores?
Answer by swincher4391(1107)

Call the square numbers x,y,z.
We know that and that x and y are both three digit numbers and that the digits are the same.
The only numbers with this property is 13^2 = 169 and 14^2 = 196. You can either know this or list out all the three digit squares until you realize it.
We add 169+196 = 365
590 - 365 = 225 <--- which is a perfect square (15^2)
So the numbers are 169,196,225.